# Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning

### v1

@@ 3.2 Caption Reward Optimization
is inspired by GRPO (Group Relative Policy Optimization) [42], which is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM $M_{V}$ performing visual captioning. For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ caption rollouts, i.e., $\mathbf{c}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cap}}(\mathbf{q}))$. Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$. The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{c}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cap}}(\mathbf{q}))}$
@@
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (1)
However, different from the math reasoning problem traditionally studied by GRPO, there are no extractable answers in the generated captions. Therefore, the widely adopted rule-based reward is no longer applicable to our problem. To tackle this, we propose to combine the rule-based reward function $r(\cdot,\cdot)$ and a reasoning LLM $M_{R}$ for reward computation. Specifically, the reward $R_{i}$ of a caption rollout $\textbf{c}_{i}$ is determined by whether a reasoner LLM $M_{R}$ could arrive at the ground-truth answer using this caption $\textbf{c}_{i}$ as contexts, which can be expressed as:
$R_{i}=r(y_{\text{gt}},y_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(y_{i})),\quad\text{where}\ y_{i}=M_{R}(P_{\text{reason}}^{\prime}(\textbf{q},\textbf{c}_{i})),$ (2)
where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise. Note that the reasoning process during reward computation is different from that in Sec. 3.1 as the tentative response is not provided here (check the ablations on this configuration in Table 3). In summary, the proposed Caption Reward Optimization offers the following benefits:
  • More accurate and task-relevant captions: As defined in Eq. (2), the reward encourages captions that facilitates successful downstream reasoning. Consequently, the MLLM learns to generate captions with less visual hallucination and more task-relevant visual details.
  • Alignment scalability to any reasoners: Once the MLLM undergoes the Caption Reward Optimization process, even LLMs that are not involved during training can be incorporated into the reasoning framework because the captions are natural languages that generalize between different LLMs. This eliminates repeated alignments as done in the traditional MLLM instruction tuning approaches. As demonstrated in Figure. 1, our RACRO reveals superior scalability during inference time, even though not been specifically trained for.
@@
Figure 4: Caption Reward Optimization reinforces captions that induce correct reasoning results via reinforcement learning with verifiable rewards (RLVR).
4 Experiments
4.1 Main Results
Evaluation.
Following VisualPRM [49], we evaluate our model on a diverse set of multi-modal reasoning benchmarks, including the MathVista (testmini) [36], MathVision (test) [48], MathVerse (vision-only) [56], MMMU (val) [55], WeMath [39], DynaMath [60], and the LogicVista [52]. We report the worst-case accuracy for DynaMath and the strict accuracy for WeMath, while for the other benchmarks, we report the overall accuracy. All evaluations are conducted using VLMEvalKit [14].
Table 1: Comparison on multi-modal reasoning benchmarks.
@@ 4.1 Main Results
InternVL3-2B-MPO 57.0 21.7 25.3 48.6 22.4 14.6 36.9 32.4
OpenVLThinker-7B 70.2 25.3 - - - - - -
R1-Onevision-7B 64.1 29.9 40.0 - - - - -
Adora-7B 73.5 23.0 - - - - - -
MM-Eureka-7B 73.0 26.9 - - - - - -
InternVL3-8B-MPO 71.6 29.3 39.8 62.7 37.1 25.5 44.1 44.3
InternVL3-14B-MPO 75.1 37.2 44.4 67.1 43.0 31.3 51.2 49.9
MM-Eureka-32B 74.8 34.4 - - - - - -
InternVL3-38B-MPO 75.1 34.2 48.2 70.1 48.6 35.3 58.4 52.8
LLaVA-OV-72B 67.5 30.1 39.1 56.8 - - - -
EMOVA-72B 69.9 - 50.0 59.7 - - - -
VL-Rethinker-72B 80.3 43.9 - 68.8 - - - -
@@
QVQ-72B-Preview 70.3 34.9 48.2 70.3 39.0 30.7 58.2 50.2
InternVL3-78B-MPO 79.0 43.1 51.0 72.2 46.0 35.1 55.9 54.6
@@
w/ VisualPRM-8B-v1.1 69.5 35.1 36.3 56.0 39.2 10.6 43.4 41.4
w/ RACRO (R1-7B) 64.5 (+0.0) 36.3 (+14.4) 39.3 (+10.5) 51.7 (+1.6) 31.5 (+7.3) 21.4 (+8.0) 46.5 (+6.9) 41.6 (+7.0)
w/ RACRO (Qwen3-8B) 68.6 (+4.1) 39.6 (+17.7) 43.7 (+14.9) 58.6 (+8.5) 32.3 (+8.1) 29.3 (+15.9) 51.9 (+12.3) 46.3 (+11.7)
w/ RACRO (R1-32B) 70.2 (+5.7) 41.5 (+19.6) 43.8 (+15.0) 58.4 (+8.3) 32.5 (+8.3) 30.3 (+16.9) 55.7 (+16.1) 47.5 (+12.9)
@@
w/ VisualPRM-8B-v1.1 74.6 27.1 45.0 56.1 46.3 17.8 48.5 45.1
w/ RACRO (R1-7B) 71.9 (+1.6) 38.9 (+13.1) 45.3 (+4.3) 58.7 (+1.4) 37.1 (+2.7) 27.0 (+7.6) 49.7 (+3.6) 46.9 (+4.9)
w/ RACRO (Qwen3-8B) 70.0 (-0.3) 41.5 (+15.7) 44.8 (+3.8) 60.6 (+3.3) 38.6 (+4.2) 26.5 (+7.1) 50.1 (+4.0) 47.4 (+5.4)
w/ RACRO (R1-32B) 74.4 (+4.1) 42.5 (+16.7) 46.6 (+5.6) 61.1 (+3.8) 31.0 (-3.4) 30.3 (+10.9) 55.0 (+8.9) 48.7 (+6.7)
@@
w/ VisualPRM-8B-v1.1 77.7 39.9 52.2 69.6 50.4 33.1 53.9 53.8
w/ RACRO (R1-7B) 75.4 (-1.4) 44.5 (+6.7) 52.9 (+2.8) 68.8 (-0.2) 48.3 (+5.2) 35.1 (+1.8) 54.6 (-0.4) 54.2 (+2.0)
w/ RACRO (Qwen3-8B) 75.5 (-1.3) 47.1 (+9.3) 54.9 (+4.8) 67.9 (-1.1) 50.0 (+6.9) 35.7 (+2.4) 58.2 (+3.2) 55.6 (+3.4)
w/ RACRO (R1-32B) 75.9 (-0.9) 46.7 (+8.9) 55.6 (+5.5) 69.0 (+0.0) 48.9 (+5.8) 37.1 (+3.8) 58.6 (+3.6) 56.0 (+3.8)
@@
We compare our method with the following baselines:
  • Proprietary models: GPT-4o [25], Claude-3.7-Sonnet [3], and Gemini-2.0-Flash [12] are general-purpose multi-modal models without dedicated reasoning optimization. In contrast, OpenAI o1 [26] is explicitly designed to enhance reasoning capabilities. These models represent the current proprietary state-of-the-art.
  • Open-source general-purpose models: LLaVA-OneVision-72B (LLaVA-OV-72B) [27], EMOVA-72B [6], and the Qwen2.5-VL series [5] are primarily pretrained or instruction-tuned on large-scale image-text datasets, without explicit reasoning alignment.
  • Open-source reasoning models: The InternVL3-MPO variants [59], OpenVLThinker-7B [13], R1-Onevision-7B [54], Adora-7B, MM-Eureka (7B/32B) [38], VL-Rethinker (7B/72B) [47], and QVQ-72B-Preview [40]. These models are explicitly fine-tuned for reasoning tasks using methods such as SFT, DPO, and rule-based reinforcement learning.
  • Multi-modal test-time scaling methods: We also compare our proposed method with multi-modal test-time scaling methods such as Best-of-N search (BoN) [49], which has been shown to effectively improve the reasoning accuracy of MLLMs. In particular, we consider the latest version, VisualPRM-8B-v1.1 [49]. Note that in Table 1, BoN is adopted independently with RACRO.
Implementation.
To validate the effectiveness of the proposed method, we perform RACRO on the following MLLMs: Qwen2.5-VL series (3B, 7B, and 32B). During training, we use R1-Distilled-7B (referred to interchangeably as R1-7B) as the reasoner to compute reward signals for all MLLMs. During evaluation, in addition to R1-Distilled-7B, we also employ its 32B counterpart and Qwen3-8B to assess the scalability of our method. For training data, we adopt ViRL39K [47], a curated dataset of 38,870 verifiable multi-modal question-answer pairs tailored for reinforcement learning. We implement the caption reward optimization algorithm using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$. For each query, four captions are sampled. The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$.
@@ Results.
Table 1 presents the performance of the proposed method. As shown, the average scores across seven datasets are significantly improved compared to the original MLLMs. For example, applying RACRO to Qwen2.5-VL-7B with a same-sized reasoner (R1-Distilled-7B) yields a +4.9% gain. Further improvements are observed when using a larger reasoner, such as a +6.7% increase with R1-Distilled-32B. Notably, when applying RACRO to Qwen2.5-VL-32B and evaluating with R1-Distilled-32B, we achieve state-of-the-art results on MathVision, MathVerse, and LogicVista, outperforming even larger models such as InternVL3-78B-MPO. Importantly, the Qwen3-8B and R1-32B reasoners are only used during inference time via a plug-and-play manner, since all trainings are conducted solely with the R1-7B reasoner. This demonstrates the scalability and generalization ability of our RACRO across both model sizes and (heterogeneous) reasoning LLM variants.
Compared to Best-of-8 (Bo8) search, RACRO also achieves superior results. Notably, for stronger MLLMs such as Qwen2.5-VL-32B, VisualPRM-8B-v1.1 provides only limited improvement (e.g., +1.6%). RACRO, on the other hand, demonstrates greater scalability to stronger reasoners: when paired with a same-sized reasoner as VisualPRM-8B-v1.1, it improves the performance of Qwen2.5-VL-32B by +3.4%. This improvement is achieved without any re-training, highlighting RACRO’s plug-and-play compatibility with more capable reasoners.
Table 2: Ablation study of different components of our RACRO.

### v2

@@ 3.2 Visual Perception Optimization (VPO)
Without the loss of generality, we describe VPO using the query-relevant caption (qcap) setting. VPO is inspired by Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a policy optimization algorithm originally developed for text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize is the MLLM performing visual captioning. For a given input pair $(I,q)$ from the training set $p_{\mathcal{D}}$, the old policy generates $G$ caption rollouts $\{o\sim\pi_{\theta_{\mathrm{old}}}(I,P_{\text{qcap}}(q))\}$ [^2] Here, we denote $o_{i}$ as the query-relevant captions qcap, rather than holistic captions cap or tentative solutions sol, as in Sec. 3.1. Let $R_{i}$ be the reward for the $i$th rollout. The normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ is the standard deviation of rewards within the group $R=\{R_{i}\}$ and $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$ is the baseline reward. Thus, the objective of VPO, following the formulation of GRPO, can be represented as:
$L(\theta)={}$ $\mathbb{E}_{(I,q)\sim p_{\mathcal{D}},o\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{qcap}}(q}))$
@@
$\Biggl[\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{qcap}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{qcap}}(q))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{qcap}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{qcap}}(q))},\,1-\epsilon_{\text{l}},\,1+\epsilon_{\text{h}}\Bigr)\hat{A}_{i}\Biggr)\Biggr]\textrm{,}$ (1)
which incorporates a surrogate loss clipped to $1-\epsilon_{\text{l}},1+\epsilon_{\text{h}}](\epsilon_{\text{l}}>0,\epsilon_{\text{h}}>0))$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown) to stabilize optimization.
@@
Figure 4: Visual Perception Optimization (VPO) reinforces captions that induce correct reasoning results via reinforcement learning with verifiable rewards. Here we omit caption penalty for simplicity.
Reward Design.
GRPO is often used in math reasoning problems (Shao et al., 2024; Guo et al., 2025), in which the reward is determined by a simple rule because the model’s output is the final solution itself. However, in our setting, the MLLM generates an intermediate caption, from which the reasoning solution cannot be directly extracted. To address this, for each caption rollout $o_{i}$, we prompt the reasoning LLM to generate a final answer, and the reward is determined by whether this answer matches the ground-truth. This is formalized as follows:
$\hat{R}_{i}=r(y_{\text{gt}},y_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(y_{i})),\quad\text{where}\ y_{i}=\operatorname{LLM}(P_{\text{reason}}^{\prime}(q,o_{i})),$ (2)
where $y_{i}$ is the answer produced by the LLM from caption $o_{i}$, $\mathbbm{1}(\cdot)$ is the indicator function, and $P_{\text{reason}}^{\prime}(q,o_{i})$ (shown in Figure 16) is the reasoning prompt different from that used for inference (template in Figure 12) as it omits the tentative solution. The reward function $r(\cdot,\cdot)$ compares the predicted answer with the ground-truth $y_{\text{gt}}$. We ablate this design choice in Sec. 4.2.
During training, we observe reward hacking (details in Sec. 4.2), where the MLLM directly solves the problem instead of performing captioning. Consequently, the model’s captioning ability does not improve. To address that, we impose a penalty on reward $\hat{R}_{i}$ if (i) $o_{i}$ leads to a correct answer; and (ii) $o_{i}$ does not contain a caption (determined by the policy MLLM $\pi_{\theta}$ via few-shot prompting):
$R_{i}=\hat{R}_{i}-\lambda\,\mathbbm{1}\big(\hat{R}_{i}=1\land\neg\text{hasCap}(o_{i})\big),$ (3)
where, $\lambda$ is a penalizing factor, and $\text{hasCap}(\cdot)$ checks if $o_{i}$ contains a caption (template in Figure 17).
In summary, VPO offers two primary advantages:
  • Generation of Reasoning-Aligned Captions: VPO uses the final reasoning outcome as a reward signal to optimize MLLMs, ensuring the captions are not merely descriptive but also functionally aligned for further reasoning. Check the quality improvement we demonstrate in Sec. 4.4.
  • LLM-Agnostic and Generalizable Improvement: VPO is an LLM-agnostic optimization, i.e., the optimized MLLMs communicate with the LLM reasoners via natural languages, and thus, the performance gains are generalizable across any LLMs. This enables a one-time alignment, which can be paired with any LLMs in a plug-and-play way without repeating VPO, as shown in Sec. 4.3.
In addition to captions, we further improve the quality of the tentative solution generated by the MLLM. As the tentative solution can be easily verified by a rule-based reward function, we apply GRPO on the MLLM. Details can be found in Appendix C. In the experiments, we optimize the MLLM with GRPO and VPO in a sequential manner, with VPO followed by GRPO.
Table 1: Comparison on multi-modal reasoning benchmarks with respect to average accuracies. The best results are bold, while the second best are underlined. : short for GPT-OSS-120B-A5B.
@@
$M$-Eureka-7B 73.0 27.9 46.1 54.9 34.7 22.6 48.3 43.9
InternVL3-8B 73.6 29.3 39.8 62.7 37.1 25.5 44.1 44.6
VL-Rethinker-7B 74.9 30.0 47.5 56.9 37.3 21.4 43.6 44.5
ReVisual-R1-7B 70.8 43.0 52.7 55.7 40.7 30.5 51.2 49.2
Ovis2-8B 71.8 25.9 42.3 57.4 27.2 20.4 39.4 40.6
InternVL3-14B 75.1 37.2 44.4 67.1 43.0 31.3 51.2 49.9
Ovis2-16B 73.7 30.1 45.8 60.7 45.0 26.3 47.4 47.0
Gemma-3-12B 56.1 30.3 21.1 55.2 33.6 20.8 41.2 36.9
$M$-Eureka-32B 74.7 36.6 51.5 62.0 37.1 33.5 58.2 50.5
InternVL3-38B 75.1 34.2 48.2 70.1 48.6 35.3 58.4 52.8
Ovis2-34B 76.1 31.9 50.1 66.7 51.9 27.5 49.9 50.6
Gemma-3-27B 59.8 39.8 34.0 64.9 37.9 28.5 47.3 44.6
QVQ-72B-Preview 70.3 34.9 48.2 70.3 39.0 30.7 58.2 50.2
@@
VL-Rethinker-72B 78.2 42.2 54.6 67.2 49.2 34.9 56.6 54.7
InternVL3-78B 79.0 43.1 51.0 72.2 46.0 35.1 55.9 54.6
Prior Caption-then-Reason Methods
ECSO 64.6 42.7 42.8 61.4 38.4 25.0 39.4 44.9
OmniCaptioner 67.5 43.3 48.0 62.2 38.7 30.5 56.2 49.5
Qwen2.5-VL series and our RAPID-enhanced counterparts
@@
w/ RAPID (Qwen3-8B) 69.6 40.8 48.6 60.9 39.1 29.3 56.4 49.2
@@
w/ RAPID (Qwen3-8B) 76.1 43.7 52.2 64.7 45.4 32.7 57.7 53.2
@@
w/ RAPID (Qwen3-8B) 76.8 47.0 54.4 67.8 48.5 36.5 60.4 55.9
w/ RAPID (GPT-A5B) 75.9 52.1 54.3 69.8 50.8 38.3 60.4 57.4
Qwen2.5-VL-72B 74.2 39.3 47.3 68.2 49.1 35.9 55.7 52.8
w/ RAPID (GPT-A5B) 75.1 53.4 56.2 72.4 52.1 37.9 59.1 58.0
4 Experiments
4.1 Main Results
@@
We compare our method with the following baselines: (i) Proprietary models, including Claude-3.7-Sonnet (Anthropic, 2025), Gemini-2.0-Flash (DeepMind, 2025) and GPT-4o (Hurst et al., 2024); (ii) Open-source general-purpose MLLMs, including Qwen2.5-VL (3B/7B/32B/72B) (Bai et al., 2025), InternVL3 (8B/14B/38B/78B) (Zhu et al., 2025), Gemma-3 (12B/27B) (Team et al., 2025a) and Ovis2 (8B/16B/34B) (Lu et al., 2024b); (iii) Open-source MLLMs specialized for reasoning, including MM-Eureka (7B/32B) (Meng et al., 2025), VL-Rethinker (7B/72B) (Wang et al., 2025a), QVQ-72B-Preview (Qwen, 2024) and ReVisual-R1-7B (Chen et al., 2025). (iv) Latest Caption-then-Reason pipelines, such as ECSO (Gou et al., 2024) and OmniCaptioner (Lu et al., 2025). For fair comparisons, we use the GRPO-optimized MLLM (Qwen2.5-VL-7B) as captioner and Qwen3-8B as the reasoner.
Evaluation
is conducted on a diverse set of multi-modal reasoning benchmarks, e.g., MathVista (testmini) (Lu et al., 2024a), MathVision (test) (Wang et al., 2024b), MathVerse (vision-only) (Zhang et al., 2024), MMMU (val) (Yue et al., 2024), WeMath (Qiao et al., 2024), DynaMath (Zou et al., 2024), and LogicVista (Xiao et al., 2024). As in recent works  (Wang et al., 2025b; Zhu et al., 2025), we use VLMEvalKit (Duan et al., 2024) for evaluation, and report the worst-case accuracy for DynaMath, the strict accuracy for WeMath, and overall accuracy for the other benchmarks.
Implementation Details of RAPID.
We perform RAPID upon the Qwen2.5-VL series (3B, 7B, 32B, and 72B) MLLMs. During training, we use R1-Distilled-7B (R1-7B) as the reasoner to compute reward signals for all MLLMs. During evaluation, we adopt Qwen3-8B  [^3] We do not use R1-7B as we found the similar-sized Qwen3-8B performs better in Sec. 4.3. (Yang et al., 2025a) and GPT-OSS-120B (Agarwal et al., 2025) as the LLM reasoners. For training data, we adopt ViRL39K (Wang et al., 2025a), a curated dataset of 38,870 verifiable multi-modal question-answer pairs tailored for multi-modal reasoning. We implement GRPO and VPO with verl (Sheng et al., 2025) with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$.
Implementation Details of GRPO.
We set the number of rollouts to 8 for the 3B/7B MLLMs and 4 for the 32B/72B MLLMs. Following Yu et al. (2025), we remove KL regularization and use the "Clip-Higher" strategy, setting $\epsilon_{\text{l}}$ to $0.2$ and $\epsilon_{\text{h}}$ to $0.25$. When reporting performance with GPRO but without VPO (e.g., and in Table 2 or the baselines in Table 3), we select the best-performing checkpoints (with perception-reasoning decouple applied) at 400, 300, and 100 steps for the 3B, 7B, and 72B MLLMs, respectively, based on the average accuracies across the seven reasoning datasets (evaluated every 50 steps). GRPO is not applied to the 32B variant, as it has already been RL-tuned.
Implementation Details of VPO.
We set the number of rollouts to 4, KL-penalty coefficient $\beta$ to $\mathrm{e}^{-3}$, and penalizing constant $\lambda$ in Eq. (3) to $0.1$. VPO is applied to the MLLM following 200 steps of GRPO [^4] For the 3B model, we observe that training with VPO after GRPO results in slight forgetting of reasoning. To mitigate this, we switch back to GRPO optimization for an additional 100 steps after VPO. (except for the 32B model, which we directly use the original model). Similar to GRPO, we select the best checkpoints at 200, 150, 100, and 100 for the 3B, 7B 32B, and 72B models according to their average accuracies on the reasoning datasets.
@@ Results.
Table 1 compares the performance of RAPID and baseline MLLMs on seven multi-modal reasoning datasets. It highlights two merits of RAPID: (i) It achieves significant performance gains on the reasoning tasks compared to the original MLLMs. For example, applying RAPID to Qwen2.5-VL-7B with a similar-sized LLM (Qwen3-8B) yields an average accuracy of $53.2\%$ (+11.2% compared to the original MLLM). Notably, when applying RAPID to Qwen2.5-VL-72B with GPT-OSS-120B as the LLM, we achieve the best average accuracy of $58\%$ across all the models compared, even surpassing proprietary MLLMs. (ii) RAPID achieves better performance-size trade-off. For example, Qwen2.5-VL-7B with RAPID (Qwen3-8B) with a total size of 15B outperforms larger models such as MM-Eureka-32B, InternVL3-38B and Ovis2-34B. Similarly, Qwen2.5-VL-32B with RAPID (Qwen3-8B) surpasses VL-Rethinker-72B and InternVL3-78B. Check Figure 1 for a better visualization of the performance-size trade-off. More analysis on the training and inference compute efficiency is provided in Appendix D. (iii) RAPID surpasses latest caption-then-reason methods (Gou et al., 2024; Lu et al., 2025), mainly due to the usage of tentative responses and VPO.
Figure 5: General benchmark Results.
Evaluation on General Benchmarks
Although RAPID is specifically designed for multi-modal math and science reasoning, we verify that it does not hurt general abilities. We evaluated the VPO/GRPO-optimized Qwen2.5-VL-7B on general benchmarks in a “non-thinking” mode, per common protocols (Yang et al., 2025a; Wang et al., 2025c). As shown in Figure 5, its performance remains on par with the original model. This confirms that our method is a targeted enhancement for reasoning that preserves the model’s general abilities. (Benchmark details and Qwen2.5-VL-3B results are in Appendix E.1).
Table 2: Ablation study of different components of RAPID (with Qwen2.5-VL-7B by default). VPO: VPO without the caption penalty; : using cap+sol for reasoning-perception decoupling.